### v2

@@ 6.1 Reproducibility of Deep Learning Bugs
6.2 Improving the Reproducibility of Deep Learning Bugs using LLMs
We conduct experiments using the LLaMA3, a 70B parameter LLM, to evaluate its effectiveness in assisting developers with the reproduction of deep learning bugs. To conduct these experiments, we construct a dataset of 40 bugs. We used stratified random sampling and collect 40 of our manually reproduced bugs, with the following distribution: 10 Training, 10 API, 10 Tensor and 10 Model Bugs. The prompt for the LLaMA3 model contained four key inputs: the bug report describing the issue, relevant code snippets providing context, the instruction to generate a code snippet that could reproduce the reported bug, and, crucially, the edit actions and component information derived from our findings. The prompt for the experiment can be found in our replication package replicationPackage . Once the model generates the code snippets, we check their ability to reproduce the bugs through manual execution and comparison with the erroneous behaviour outlined in the bug reports. Furthermore, we compare the generated code with the ground truth code snippets curated during our manual bug reproduction process.
The results of our experiments are as follows:
6.2.1 Statistical Significance Tests for our Experiments
To assess the statistical significance of the LLM results, we employ McNemar’s test pembury2020effective , which is designed to detect differences between paired samples of two binary variables. The null hypothesis of McNemar’s test is that the LLM’s bug reproduction ability is the same with or without hints. We organize the samples in a 2x2 contingency table, as shown in Tab. 14. We explain the values present in the table in detail below.
  • 23: This value indicates the number of bugs that were reproduced with/without the hints.
  • 2: This value indicates the number of bugs that were reproduced in the absence of the hints, but could not be reproduced in the presence of the hints.
  • 10: This value indicates the number of bugs that were reproduced in the presence of the hints, but could not be reproduced in the absence of the hints.
  • 5: This value indicates the number of bugs that could not be reproduced with/without the hints.
Table 14: McNemar’s test contingency table for bug reproduction with and without augmentation
Using Hints: Using Hints: Total
Reproducible Not Reproducible
Without Hints: Reproducible 23 2 25
Without Hints: Not Reproducible 10 5 15
Total 33 7 40
It is important to note that the LLM may struggle to reproduce some bugs without full code context, even with our augmentation. Despite these limitations, our results, when compared to the LLM’s performance without augmentation, yield a McNemar’s test p-value of 0.043, which is less than 0.05. The results provide strong evidence to reject the null hypothesis and accept the alternate hypothesis, indicating that the LLM has a statistically significant improved ability to reproduce bugs when provided with hints compared to without hints.
The recommendations of LLMs are stochastic by design, and different runs with the same prompt can produce different reproducible snippets. This introduces an additional source of variability which may invalidate our previous observations. To address this issue, we repeat the code generation task for each bug five times and record their rates of success in bug reproduction. Since the bug reproduction rates are continuous percentage values rather than binary outcomes, the McNemar test cannot be used. To overcome this limitation, we use the RM-ANOVA test rutherford2011anova . RM-ANOVA is suitable for this scenario because we have a within-subjects design with one independent variable (augmentation with hints or not) and one dependent variable (bug reproduction rate) measured repeatedly across trials. The test yields a significant p-value of $5.8\times 10^{-5}$, further demonstrating the positive impact of hints on bug reproduction abilities of large language models.
@@ Acknowledgements
Appendix A Flowchart for Manual Analysis

### v3

@@ 6.1 Reproducibility of Deep Learning Bugs
6.2 Challenges in Reproducing Deep Learning Bugs: A Comparison between Stack Overflow and GitHub
Reproducing deep learning bugs from Stack Overflow posts and GitHub repositories presents different challenges and requires varying levels of effort. The main differences lie in the availability of code and data, environmental setup, context, completeness, and reproducibility expectations. We explain them briefly as follows.
First, GitHub repositories often provide the complete codebase and sometimes the accompanying datasets, whereas Stack Overflow posts typically include small code snippets without the broader context and data. Hence, we need more effort in reconstructing the code and generating synthetic data when reproducing bugs from Stack Overflow. Second, GitHub repositories commonly include setup instructions, dependency files, and development environment configurations, making it easier to recreate the original environment accurately. On the contrary, Stack Overflow posts rarely provide such details, requiring guesswork about software versions, dependencies, and environment settings. Finally, GitHub issues tend to have more contextual information, such as detailed problem descriptions, steps to reproduce a bug, and error logs supporting root cause analysis of the issue. Stack Overflow posts may lack this level of detail, making it harder to comprehend and reproduce a bug accurately.
Despite the additional resources available in GitHub repositories, reproducing bugs from them still presents several challenges:
  • Incomplete or outdated repositories: Critical files, dependencies, or code changes may be missing or outdated, making it difficult to reproduce the exact environment and conditions under which a bug occurred.
  • Large and complex codebases: Deep learning projects often have extensive and intricate codebases, requiring significant time and expertise to set up and navigate them.
  • Proprietary or sensitive data: Many projects may involve proprietary data that cannot be shared publicly, making it challenging to reproduce data-related bugs without access to the original data.
  • Insufficient documentation: Many GitHub repositories fail to provide clear documentation on project architecture, installation steps, and expected behaviours. The lack of clear documentation makes it difficult for developers to reproduce bugs, as they face challenges in understanding the codebase and configuring the project locally.
  • Dependency management: Resolving dependency conflicts or managing compatibility issues across different project versions can be a significant challenge when reproducing bugs.
To address these challenges, we recommend several changes or adaptations to the status quo as follows. First, Stack Overflow should leverage its user base and search functionality for discussing common deep-learning bugs and quick solutions, while GitHub should be the primary platform for in-depth bug reproduction and resolution. Second, Stack Overflow should introduce a specialized format for deep learning bug reports, expanding on its traditional minimum working example approach to include dataset details, framework versions, hardware specifications, and reproducible code snippets. GitHub repositories should be kept up-to-date, well-documented, and complete. Finally, both platforms should incentivize active participation in bug reporting and resolution through specialized badges, reputation points, or recognition. Stack Overflow could introduce “Deep Learning Debug Rooms” for real-time, collaborative problem-solving, and GitHub could highlight top contributors in bug resolution. By combining Stack Overflow’s community-driven approach with GitHub’s comprehensive project management features, we can develop a more efficient ecosystem for deep learning bug resolution, leading to more robust deep learning systems.
@@ Acknowledgements
Appendix A: Flowchart for Manual Analysis
@@
Appendix B: Improving the Reproducibility of Deep Learning Bugs using LLMs
We conduct experiments using the LLaMA3, a 70B parameter LLM, to evaluate its effectiveness in assisting developers with the reproduction of deep learning bugs. To conduct these experiments, we construct a dataset of 40 bugs. We used stratified random sampling and collect 40 of our manually reproduced bugs, with the following distribution: 10 Training, 10 API, 10 Tensor and 10 Model Bugs. For each of these four bug categories, we ensure that the dataset contains an equal number of easy and difficult bugs, determined by the number of edit actions required to reproduce the bug. Specifically, within each bug category, there are 5 bugs that require a relatively small number of edit actions (classified as easy), and 5 bugs that require a larger number of edit actions (classified as difficult). The prompt for the LLaMA3 model contained four key inputs: the bug report describing the issue, relevant code snippets providing context, the instruction to generate a code snippet that could reproduce the reported bug, and, crucially, the edit actions and component information derived from our findings. Fig. 6 demonstrates an example prompt for our experiments. Once the model generates the code snippets, we check their ability to reproduce the bugs through manual execution and comparison with the erroneous behaviour outlined in the bug reports. Furthermore, we compare the generated code with the ground truth code snippets curated during our manual bug reproduction process.
Figure 6: Prompt for LLaMA3
The results of our experiments are as follows:
Statistical Significance Tests for our Experiments
Since the recommendations of LLMs are stochastic by design, different runs with the same prompt can produce different reproducible snippets. This introduces an additional source of variability which may invalidate our previous observations. To address this issue, we repeat the code generation task for each bug five times and record their rates of success in bug reproduction. To assess the statistical significance of the LLM results, we employ the RM-ANOVA test rutherford2011anova . RM-ANOVA is suitable for this scenario because we have a within-subjects design with one independent variable (augmentation with hints or not) and one dependent variable (bug reproduction rate) measured repeatedly across trials. The test yields a significant p-value of $5.8\times 10^{-5}$, demonstrating the positive impact of hints on bug reproduction abilities of large language models. Additionally, we have used the partial-eta squared ($\eta_{p}^{2}$) metric for the effect size, as it is a common measure in ANOVA representing the proportion of total variance attributable to a factor adjusted for other factors in the model. The $\eta_{p}^{2}$ for the Guidance factor (hint augmentation) is 0.188324, which is considered a large effect size. This substantial effect size, coupled with the significant p-value, provides strong evidence for both the statistical and practical significance of using hints to improve LLM performance in bug reproduction tasks.